TOPIC 9: Domain Range of a Rational functions Find the domain and range of the following rational function

1.f(x)=2 x+1

2.f(x)=3x x+3

3.f(x)=3-x x-7

4.f(x)=2+x x

5.f(x)=(x+1) x^2-1


1
Expert's answer
2021-10-11T05:32:26-0400

1) "f(x)=\frac{2}{x+1}. Domain: (-\infin,-1)\cup(-1,\infin).Range:(-\infin,0)\cup(0,\infin)"


2) "f(x)=\frac{3x}{x+3}. Domain: (-\infin,-3)\cup(-3,\infin).Range:(-\infin,3)\cup(3,\infin)"


3) "f(x)=\frac{2}{x+1}. Domain: (-\infin,-1)\cup(-1,\infin).Range:(-\infin,0)\cup(0,\infin)"


4) "f(x)=\frac{2+x}{x}. Domain: (-\infin,0)\cup(0,\infin).Range:(-\infin,1)\cup(1,\infin)"


5) "f(x)=\frac{x+1}{x^2-1}. Domain: (-\infin,1)\cup(1,\infin).Range:(-\infin,0)\cup(0,\infin)"